### v1

# 2]organization=European Synchrotron Radiation Facility, addressline=71 Avenue des Martyrs, city=Grenoble, postcode=F-38000, state=France, country=

Abstract
A diamond superlattice with a period length of 3.54 $\mu$m was grown on a high quality straight (100) diamond plate with the method of Chemical Vapour Deposition (CVD). A sinusoidal varying boron doping profile resulted in a periodic variation of the lattice constant, and in turn four sinusoidally deformed (110) planes with a period length of 5.0 $\mu$m and an amplitude of 0.138 nm. A channeling experiment was performed with the 855 MeV electron beam of the Mainz Microtron MAMI accelerator facility. Part of the impinging electrons perform sinusoidal oscillations resulting in the emission of quasi-monochromatic $\gamma$ radiation. A clear peak was observed with a large sodium iodide scintillation detector close to the expected photon energy of 1.33 MeV. Gross properties like photon energy, width and intensity of the peak can be reproduced fairly well by idealized Monte-Carlo simulation calculations. Based on the latter, prospects of applying such $\gamma$ radiation sources are addressed with the example of the photonuclear reaction 100Mo($\gamma$,n)99Mo at 14.3 MeV to produce the medical important 99mTc isotope.
keywords:
[…]
Introduction
Intense narrow-band multi-MeV photon beams have a large potential in fields of basic and nuclear research, medicine and technical applications [1, 2, 3, 4]. Having in mind in particular photonuclear reactions for application purposes, the production mechanisms of such beams should be targeted, from an economical point of view, on electron accelerators with moderate energies in the GeV range or below. Bremsstrahlung radiation produced by an electron beam incident on a high-Z target has been widely used for many decades. Alternatively, also channeling radiation has been considered as an intense radiation source as well [5]. However, both sources feature broad band radiation characteristics. Facilities based on laser Compton scattering are already operational or currently under construction [3, Table 10] by which nearly mono-energetic $\gamma$-ray beams of high intensity can be produced.
Figure 1: Effect of a periodic variation of the lattice constant. A superlattice (undulator) crystal originates as shown by the red colored full line for one period. The effect of the tetragonal distortion of the equilibrium cubic unit cell has been described by M.B.H. Breese [6]. The electron beam enters in $z$ direction, called forward direction, which coincides with the [110] direction of the plane diamond backing (host) crystal. The $z^{\prime}$ direction is called undulator direction. The two coordinate systems $(x,z)$ and the $(x^{\prime},z^{\prime})$ are interconnected by a rotation transform with an angle $\delta$. Planes are formed perpendicular to the plane of drawing. Figure taken from [7].
Figure 2: Transverse potential landscape of a boron doped (110) diamond superlattice. Shown is one period with length of 5.0 $\mu$m, centrifugal potential omitted. A full sinusoidal period starts at $x=z$ = 0. Notice, the longitudinal $z$ coordinate is squeezed by four orders of magnitudes compared with the transverse $x$ one. It is shown in Methods Monte-Carlo simulation calculations that the barrier heights are modified as functions of $z$ by centrifugal forces. The quantity $\overline{x}$ is defined by Eq. (4).
In this work an alternative route towards narrow band photon sources was elaborated which is based on the channeling process of ultra-relativistic electrons in periodical bent crystals. The scheme resembles magnetic undulators, however, with period lengths in the multi-$\mu$m range rather than on a centimeter scale. Such undulator-like radiation sources in the MeV range and beyond have been comprehensively discussed by Korol and Solov’yov, see e.g. [8] and references cited therein.
The experimental challenge is to fabricate crystalline undulators with such small period lengths. First attempts were undertaken with strained layer Si1-xGex undulator crystals grown at the Aarhus University [9]. Of greater interest are undulators based on low $Z$ = 6 diamond, because of its radiation hardness, which can be produced by Chemical Vapour Deposition (CVD) with periodically varying concentration of boron. It is based on the functional dependence of the lattice constant on the boron density, e.g. [10, Fig. 1]. As schematically shown by Fig. 1, oscillating (110) planes result with a potential landscape depicted in Fig. 2.
An electron entering the crystal at the coordinates $x=z$ = 0 nearly parallel to the $z$ axis may be captured in the potential pocket with a barrier hight of 24 eV, and performs oscillations, thereby following the bent sinusoidal shape of the (110) plane. By means of the accompanied acceleration channeling radiation and undulator-like radiation will be emitted, respectively. The radiation is strongly forward directed, for undulator-like radiation typically in the order of 0.5 mrad for 1 GeV electrons. The emitted photon energy is in the $\gamma$-ray domain, for a period length of a few $\mu$m typically in the MeV range. It scales with the electron beam energy squared and inversely with the undulator period length. The capture probability and in turn the intensity of the $\gamma$ radiation requires high quality electron beams with a low angular spread. The reason is that the transverse energy $pv/2\cdot\vartheta_{x}^{2}$ is the relevant quantity for the dynamics of the electron in a potential landscape as shown in Fig. 2. The quantity $pv$ is in essence the electron beam energy and $\vartheta_{x}$ the angle of the electron with respect to the $z$ axis. Collisions of the beam electron with atomic nuclei and electrons comprising the crystal, potentially result in an increase of the transverse energy and de-channeling processes. That is another reason why the low $Z$ = 6 material of diamond is a good choice.
An experiment was performed by us at planar (110) channeling of electrons in a four-period diamond undulator with a period length $\lambda$ = 5.0 $\mu$m. Of paramount importance was the successful fabrication by the CVD process and diagnostics of such a boron doped superlattice. The experiment has been conducted with the high quality 855 MeV electron beam of the Mainz Microtron MAMI accelerator facility. Of crucial importance for the successful peak search at the predicted energy of 1.33 MeV was a strategy developed prior to the experiment. To this end Monte-Carlo simulation calculations were performed with a computer code developed by us, which facilitated the experimental search with optimized parameter predictions.
Our computer code fairly well reproduces the basic measured characteristics of the quasi-monochromatic $\gamma$ radiation as the peak energy, its width and the intensity. This fact encouraged us to predict, as an example, features of intense quasi-monochromatic 14.3 MeV radiation from an eight-period diamond undulator with $\lambda$ = 5 $\mu$m and a 3 GeV electron beam. Such a $\gamma$-ray source may be of potential interest for inducing nuclear ($\gamma,n$) reactions.
Figure 3: Measured boron doping profile. A 250 x 250 µm² crater allowed for the analysis of the material as a function of depth with ToF-SIMS. On the left scale the boron concentration is shown, on the right one the corresponding relative lattice expansion $\Delta$a/a, according to the "atomic volume interpolation" [10], in the following called modified Vergard’s law.
Figure 4: Depth profile taken with the Rocking Curve Imaging (RCI) technique [11, 12]. The signal is the integrated intensity of Bragg diffraction peaks at (111) planes with 17 keV photons.
Results
Superlattice
[…]
The boron doping profile was examined with the Time-of-Flight Secondary Ion Mass Spectrometry (ToF-SIMS) method. Figure 3 shows the depth profile. The SIMS data indicate that the growth process reproduce the four sinusoidal periods, irrespective of the analyzed region. As designed, the layers are consistent with a period length of  3.54 $\mu$m along the [100] growth direction. A slight variation of the doping maxima and minima from one period to the other can be observed for the layers apart from the substrate.
Fig. 4 depicts an enlarged map of the integrated intensity of the Bragg peak, taken with the Rocking Curve Imaging (RCI) technique [11, 12] at the European Synchrotron Radiation Facility, Grenoble. The depth profile in horizontal direction is shown by the vertical strips. The structure corresponds to the doping profile of Fig. 3 when interpreting low intensity (dark blue) with the region of large boron doping. In vertical direction the image shows that the layers are very homogeneous. This result has been confirmed over more than 1 mm along the crystal. A detailed quantitative analysis is in progress.
Figure 5: Experimental setup at MAMI, not to scale. Radiation can be observed at virtual collinear geometry of the electron beam direction and the sodium iodide NaI(Tl) radiation detector. The inset depicts the undulator crystal with the position of three SIMS craters.
Accelerator configuration
The experimental setup at the Mainz Microtron MAMI is shown at Fig. 5. The electron beam is focused onto the diamond crystal in the target chamber with a goniometer, and thereafter deflected by a dipole magnet into the beam dump. The target can be rotated in all 3 spatial directions with an accuracy of 35 $\mu$rad. The target is aligned using the signal of the ionization chamber that registers the scattered electrons. The orientation of the electron beam parallel to planes increases the scattering probability and, in turn, the signal of the ionization chamber. This method allows rapid alignment to a crystal plane.
The radiation in the forward direction is detected with a 10 inch sodium iodide crystal, NaI(Tl), positioned at a distance of about 8.5 m from the target. A movable cylindrical tungsten aperture defines the observation direction. It has a length of 261 mm and a bore of 4 mm diameter which was reduced with an additional 28.8 mm long tungsten cylinder inset with an inner diameter of 2 mm. The alignment of the aperture was done by maximizing the detected bremsstrahlung from an aluminum foil.
Figure 6: (a) Experimental spectra of the undulator chip per electron, including the 194 $\mu$m thick backing crystal. Spectra taken with the NaI(Tl) detector. Spectrum in green colour for de-tuning the aperture from forward direction by $\Delta\theta_{x}$ = - 0.2 mrad, and in blue colour by + 0.2 mrad. Spectra taken at typical beam current of 3.95 pA. (b) Low energy part of difference spectrum. Background spectrum taken at $\theta_{x}$ = +0.2 mrad aperture position (blue) subtracted from spectrum taken at -0.2 mrad (green). (c) and (d) Monte-Carlo simulated spectra per electron without a backing crystal. Ideal sinusoidal boron doping profile and infinitesimal small aperture size assumed. Mean of 200 single simulations.
Spectra
The peak search was performed at an electron energy of 855 MeV. The angle of the tungsten aperture was varied around the forward direction. There exist two favourable observation angles. One is the direct forward direction $\theta_{x}=0$ for which the radiation direction defining aperture is on-axis with the electron beam direction, i.e., in Fig. 1 with the $z$ axis. The other one is the undulator radiation direction in which the aperture is de-tuned by the angle $\theta_{x}=\delta$ = - 0.174 mrad into the $z^{\prime}$ direction. For both directions $\theta_{y}=0$ was chosen.
The peak energy of the undulator radiation to be sought for can be estimated with the well known relation
$\hbar\omega=k\frac{4\pi\gamma^{2}\hbar c}{\lambda\left(1+K^{2}/2+\gamma^{2}(\delta-\theta_{x})^{2}\right)}$ (1)
with Lorentz factor $\gamma$ = 1674.2, the reduced Planck constant $\hbar$, and $c$ the speed of light. In lowest order $k$ = 1 a peak energy $\hbar\omega$ = 1.33 MeV results at observation in undulator forward direction $\theta_{x}$ = $\delta$, $A_{U}$ = 0.138 nm, $\lambda$ = 5.0 $\mu$m, and undulator parameter $K=\gamma\cdot A_{U}\cdot 2\pi/\lambda$ = 0.291.
Fig. 6 depicts spectra taken with the sodium iodide detector (a) and (b), as well as results of corresponding simulation calculations (c) and (d).
Discussion
Spectra
The spectrum in green color of Fig. 6 exhibits a clear narrow peak with an energy of 1.30 MeV. It appears only at a negative observation angle $\theta_{x}$ = - 0.2 mrad, while at the positive angle of + 0.2 mrad mainly background is observed, see spectrum in blue colour. The result shows that our method for a peak search, as proposed in [7, Fig. 11] and applied for an unsuccessful experiment, is rather sensitive and enables the creation of a difference spectrum shown in Fig. 6 (b). In contrast to the above mentioned Si1-xGex strained layer crystal, the diamond undulator structure cannot be separated from the host diamond crystal. As a consequence, one is always faced with radiation from the backing, the thickness of which can be minimized to values in the several $\mu$m range. However, in our experiment the thickness in [100] direction was 194 $\mu$m, an actually rather large value in comparison with the undulator layer - actually a factor of 14 thicker. Therefore, a huge channeling radiation contribution is present, and consequently the difference spectrum is afflicted with large uncertainties, in particular above about 2 MeV.
Figs. 6 (c) and (d) depict Monte-Carlo simulated spectra under idealized conditions. The calculations are based on the formalism of [7], extended to simulate also sinusoidal profiles. Details are described in Methods Monte-Carlo simulation calculations. With some care in mind, the simulated spectrum (d) will be compared with the experimental one (b). The simulation reproduces with 1.29 MeV well the measured peak energy of 1.30 MeV. However, the experimental width is broader, and the total intensity significantly lower. The Bragg diffraction image of Fig. 4 may indicate that the amplitude decreases by relaxation of the stress, a fact not taken into account in the simulation calculation. In addition, an infinitesimal small aperture was assumed which is also unrealistic, not to mention that also the simulation model has systematic errors.
Finally we mention that the peak should also be observable by aligning the crystal into the (111) plane for which an enhanced peak signal is expected, since the transverse window in which the electron is captured for channeling may increase in comparison with the (110) alignment.
Prospects
Based on our observation in combination with the Monte-Carlo simulation result, we address the question of applications of narrow-band $\gamma$ radiation emitted from a boron-doped diamond superlattice. Main restrictions originate from the superlattice structure, rather than the electron accelerator facilities for which technologies are well developed. The basic underlying sensitive parameter is the relaxation length $L_{r}$ of the undulating structure as function of period length $\lambda$, number of periods $N_{U}$, amplitude $A_{U}$ and doping profile. It is probable a rather involved quest of material science which must be tackled experimentally, perhaps assisted by simulation calculations.
In the following we make the idealized assumption of an undulator with $L_{r}=\infty$, $\lambda$ = 5.0 $\mu$m, $N_{U}$ = 8, and $A_{U}$ = 0.092 nm. The boron doping profile was assumed to be triangular with a maximum boron concentration $c_{b,max}$ = 6.0 $\cdot 10^{20}$/cm3 resulting in a constant radius of curvature of 10.8 mm. In contrast to a sinusoidal doping profile there are no regions with infinite bending radius and vanishing acceleration. As a consequence, the radiation emission is maximized. An electron beam with an energy of 3 GeV was assumed. These parameters were identified from Monte-Carlo simulation calculations to produce a secondary narrow-band $\gamma$ radiation beam well suited for applications with photonuclear ($\gamma$,n) processes. The chosen parameter set does not originate from any optimization processes.
Figure 7: Simulated photon spectrum per electron. Beam energy 3 GeV, assumption of an ideal triangular boron doping profile, parameters as specified in the text. An infinitesimal small aperture was assumed, positioned at $\theta_{x}=\delta=-0.116$ mrad, and $\theta_{y}=0$. The narrow structures at 30 MeV and 46 MeV are interpreted as the second and third harmonic of the dominating peak at 14.3 MeV. A mean of 200 single simulations is shown.
[…]
The following crude estimates will be performed for the medical important 100Mo($\gamma$,n)99Mo reaction, as an example, with 10 enriched 100Mo disks of 3 mm diameter and 1 mm thickness each as secondary target, one after another. The total mass is 0.7 g. With the reaction cross section of 0.15 b [2, Fig. 4] and an irradiation time to reach half the saturation activity one obtains a specific activity of about 4.5 GBq of 99Mo (T1/2 = 65.9 h).
The heat load at a 100 $\mu$m thick diamond target plate at a beam current of 100 $\mu$A can be estimated to be about 10 W which seems to be manageable. The total power of the electron beam amounts to 300 kW. If energy recovery would be feasible, the power consumption and the radioactive waste in the beam dump can be reduced.
Although the proposed generator with the assumed parameters may be less intense in comparison with existing and proposed ones [13, 14], a detailed feasibility study with optimized undulator parameters and electron beam energy seems to be worth the effort. It should include considerations on the high energy channeling background, the production of other medical interesting radionuclides, as well as considerations on the radiation damage of the diamond undulator crystal.
Methods
Superlattice design and fabrication
The design profile of the boron doping in [100] direction is of sinusoidal shape with the minimum and maximum boron content $x_{b,min}$ = 5.67 $\cdot 10^{-4}$ and $x_{b,max}$ = 56.7 $\cdot 10^{-4}$, corresponding to a boron concentration $c_{b,min}$ = 1.0 $\cdot 10^{20}$/cm3 and $c_{b,max}$ = 10.0 $\cdot 10^{20}$/cm3, respectively. Notice, a boron content is defined here as $x_{b}=n_{b}/n_{C}$ with $n_{b}$ the boron concentration in units cm-3 and $n_{C}=1.763\cdot 10^{23}$/cm3 the density of carbon atoms for diamond. The boron content reads as function of the $z$ coordinate for the first half period $x_{b}(z)=x_{b,min}+(x_{b,max}-x_{b,min})/2\cdot(1-\cos[2\pi z/\lambda])$, $20\leq z\leq\lambda/$ 2. The continuation to $\lambda/2<z\leq\lambda$ can be obtained by symmetry considerations.
The lattice expansion for the "atomic volume interpolation" [10, Fig. 1], also called linear "Modified Vegard’s low" reads $\Delta d_{110}/d_{110}(z)=k\cdot x_{b}$ with the slope constant $k=0.0962$. A differential equation $dx_{b}/dz=-k\cdot x_{b}(z)$ must be solved to obtain the nominal bending of the (110) channel. Notice that beside the boundary condition $x_{b}(z\rightarrow 0)=0$ also $dx_{b}/dz(z\rightarrow 0)=0$ must be fulfilled. The solution is schematically shown in Fig. 1.
We note that when fabricating undulator devices attention must be paid to both the quality of the doped overlayers, and the quality of diamond substrate. Our application requires single-crystal diamond substrates with the main surface orientation (100), with a miscut angle of 2 degrees within one surface, and with highly perfect bulk and surface crystal quality. In addition, such diamond plates should be available with large areas (several mm2) free of extended defects such as dislocations. The diamond surface plays also an important role for the quality of the overgrown layers on it, therefore it must be as flat as possible. With special mechanical- chemical polishing, we reach the flat surface of less than 1nm r.m.s. Our final choice is high-crystalline-quality HPHT type IIa substrates with the size around 5x5mm2. The thickness of the diamond substrate varies from 80 to 270 µm from one side to the other, due to the miscut angle of 2 degrees. The Diffraction Rocking Curve Imaging (RCI) technique shows not many defects present in the bulk: some dislocations, no stacking faults, no growth sector boundaries, no inclusion and very few defects introduced by surface processing.
For the growth of the boron doped diamond layers, the diamond sample was first subjected to a tri-acid cleaning process consisting of 1 volume of HClO4, 4 volumes of HNO3, and 3 volumes of H2SO4 for 1 hour under boiling conditions. This step effectively removes residual graphitic surface layers. The sample was then rinsed with acetone, followed by alcohol and isopropyl alcohol (IPA). Subsequently, the sample was loaded into a microwave-assisted CVD (Chemical Vapor Deposition) reactor for diamond growth. The reactor plasma was ignited under a hydrogen atmosphere at a pressure of 44 mbar and a power of 250 W. The sample was exposed to this plasma for 1 hour to ensure uniform temperature distribution across both the sample and the reactor interfaces.
The growth process itself is a key innovation to achieve the desired doping profile required for the specific application. A tailored dilution stage, with precise control of the mass flow controllers, was employed to ensure a sinusoidal boron doping pattern in the gas phase. During the growth phase, a gas flow of 100 sccm of H2 and 4 sccm of CH4 was maintained, with a variable B2H6 concentration injected into the gas mixture. Based on established calibration curves, the expected boron doping level in the diamond ranges from 1 $\cdot 10^{20}$ cm-3 to 10 $\cdot 10^{20}$ cm-3. The pressure was kept at 44 mbar with a power setting of 250 W. Temperature control was achieved using a pyrometer, maintaining an average temperature of 850°C throughout the process. Considering the growth rates, each doping cycle lasted for 6015 seconds with the growth rate of 0.58 nm/s.
Monte-Carlo simulation calculations
We recall from Ref. [7, Eq. (6) and (11)] that the spectral power density can be calculated if the angular trajectories $\vartheta_{x,y}(z)$ are known. The simulation calculation code described in that paper holds only for a triangular doping profile resulting in a constant bending radius for each half period. However, as described in Methods Superlattice design and fabrication the superlattice was produced with a sinusoidal doping profile. This fact required some modifications of the simulation code.
Let $u(x^{\prime})$ be the planar potential in the $(x^{\prime},z^{\prime})$ coordinate system, see Fig. 1. Than a family of sinusoidal varying potentials may be described by
$u(x^{\prime},z^{\prime})=u(\mbox{mod}[x^{\prime}+A_{U}\sin(2\pi z^{\prime}/\lambda),d_{p}])$ (2)
with $d_{p}$ the planar distance. Transformation into the $(x,z)$ system by rotation with an angle $\delta$, taking into account the centrifugal potential, gives
$U(x,z)=u(\overline{x})-pv/R(z)~{}x$ (3)
$\overline{x}=\mbox{mod}[x+z\cdot\delta+A_{U}\sin(2\pi z/\lambda),d_{p}]$ (4)
$1/R(z)=(1/R_{min})\sin(2\pi z/\lambda)$ (5)
. In Eq. (4 )was approximated $\overline{z}=z$, $x^{\prime}=x$, $\sin(\delta)=\delta$, and $\cos(\delta)=1$. It is $pv$ = $\beta\gamma m_{e}c^{2}$ with $m_{e}c^{2}$ the rest mass of the electron and $\beta=v/c=\sqrt{1-1/\gamma^{2}}$ the reduced speed $v$ of the electron. The additional term describes the centrifugal potential with $R(z)$ the bending radius, which is a function of the longitudinal coordinate $z$, and $R_{min}$ its minimum value. The two functions $u(\overline{x})$ and $-(pv/R(z))~{}x$ are depicted in Fig. 2 and Fig. 8, upper part, respectively. The superimposed potential is shown in Fig. 8, lower part.
Figure 8: Upper: Centrifugal potential part $-(pv/R(z))~{}x$ of Eq. (3). As the particle proceeds in $z$ direction the potential part $u(\overline{x})$ depicted in Fig. 2 is modified for any specific $z$ by the centrifugal term. Only the positive part of the potential is shown. Lower: Total transverse potential landscape $U(x,z)$, Eq. (3). The big hills at right and left originate from the centrifugal potential. However, gradient forces originating from $-(pv/R(z))~{}x$, see upper part, act only in $x$ direction.
It is important to understand that the centrifugal potential must not be considered as a real potential from which gradient forces can be derived. The reason is that for the movement of the electron in $z$ direction a constant speed has been assumed, i.e. no forces act. Insofar the $z$ dependence of the potential must be treated as a parameter, rather than as a variable.
An electron entering the potential landscape at the window close to $x=z$ = 0 may be captured in the shown channel. Collisions may result in de-channeling and the free particle with a positive transverse energy may bounce against the right potential hill, will be reflected and may enter in regions of negative centrifugal potential below the plane $U(x,z)$ = 0. As it proceeds further in $z$ direction it may be reflected by an increasing centrifugal potential, like the left hill, in opposite direction.
At a step from $k\rightarrow k+1$, with $k$ an integer, the transverse $x$ coordinate changes from $x_{k}\rightarrow x_{k+1}=x_{k}+\Delta x_{k}$, the longitudinal one from $z_{k}\rightarrow z_{k+1}=z_{k}+\Delta z_{k}$, and the potential $U(x,z)$ from $U(x_{k},z_{k})\rightarrow U(x_{k+1},z_{k+1})$. At the simulation each step was subdivided into (i) a move $x_{k}\rightarrow x_{k+1}+\Delta x_{k}$ at constant $z_{k}$, and (ii) in a move from $z_{k}\rightarrow z_{k+1}+\Delta z_{k}$ at constant $x_{k+1}$. At move (i) potential forces -$\partial U(x,z)/\partial x$ act in transverse $x$ direction at conserved transverse energy $E_{\bot,k}$ and the longitudinal step size can be determined as
[…]
For each $\Delta z_{k}$ the scattering angle changes $\Delta\vartheta_{k}^{at}$ and $\Delta\vartheta_{k}^{el}$ by collisions with atoms and electrons, respectively, are determined randomly by a Monte Carlo simulation. The new scattering angle is given with $\Delta\vartheta_{k}^{sc}$ = $\Delta\vartheta_{k}^{at}$ + $\Delta\vartheta_{k}^{el}$ by
$\vartheta_{k+1}(z_{k})$ $=$ $\vartheta_{k}^{kin}+\Delta\vartheta_{k}^{sc}=$ (8)
=±$=\pm\sqrt{2\big{(}E_{\bot,k}-U(x_{k},z_{k})\big{)}/pv}+\Delta\vartheta_{k}^{sc}.$
The first term $\vartheta_{k}^{kin}$ is the instantaneous scattering angle associated with the kinetic energy $E_{\bot,k}-U(x_{k},z_{k})$. Notice, at move (ii) forces are assumed to be absent, meaning that the kinetic energy $T_{\bot,k+1}$ remains constant and the scattering angle does not change.
A scattering angle change $\Delta\vartheta_{k}^{sc}$ results in a change of the transverse energy
$E_{\bot,k+1}-E_{\bot,k}$ $=$ (9)
$pv/2\cdot(\Delta\vartheta_{k}^{sc})^{2}\pm\sqrt{2pv\cdot\big{(}E_{\bot,k}-U(x_{k},z_{k})\big{)}}~{}~{}\Delta\vartheta_{k}^{sc}.$
Further details including the choice of the sign are described in [15] behind Eq. (33).
Finally, the angular trajectory $\vartheta_{x}(z)$ results from the outlined iteration process. Calculation can be performed in the potential of the straight reference trajectory if the solution is transformed back into the bent system by adding the term $\delta\cdot(1-\cos(2\pi/\lambda\cdot z))$ to it.
The code runs on the Wolfram Mathematica 14.2 platform on a personal computer. It delivers useful results in moderate real times of less than about 10 minutes.
Acknowledgements
We appreciated help of Daniel Dominguez, Philipp Kompa, and Yaideny Rodriguez during measurements at MAMI.
We thank Andrei V. Korol and Andrey V. Solov’yov for fruitful discussions and organizational support of the project.
Declarations
This work has been financially supported by the European Innovation Council (EIC) Pathfinder TECHNO-CLS project 101046458 and the N-light project H2020-MSCA-RISE N-LIGHT No. 872196.
Data availability statement
All data are freely available.
Authors contributions
Boron-doped diamond growth: David Eon
Diamond preparation: Thu Nhi Tran Caliste
Crystal characterization: Patrik Straňák, Lutz Kirste, Thu Nhi Tran Caliste, Rébecca Dowek
Model development: Hartmut Backe, Werner Lauth, Thu Nhi Tran Caliste
ESRF beamtime: Rébecca Dowek, Pierre Everaere, Simon Bénichou, José Baruchel, Thu Nhi Tran Caliste
MAMI beamtime: Werner Lauth, Pascal Klag, Rébecca Dowek, Thu Nhi Tran Caliste
Editor: Hartmut Backe
[…]